1) a steel wire of lenght l has a magnetic moment M , if the wire is bent in semi-circular arc what would be the moment   ans 2M /pie
 
2) calculate the magnetic momemt at a pt on the equator at a distance 20cm from the centre of a magnet of dipole moment 5Am^2  ans :6.25 *10^ -5 wB/m^2
 
3)the magnetic induction at a pt P on the axis of a short magnet is equal to magnetic induction at  a pt Q on the equatorial line .compare the distances from the centre of the dipole  ans : 2^ 1/3  : 1
 
4) a short bar magnet of moment of 5Am^2 is placed with its axis in the magnetic meridian  and its N -pole facing south ,if the neutral pts are observed at a distace of 10cm form hte centre of the magnet , find Bh . the bar magnet is now reversed now find the neutral pts  ans :8 cm from the centre on the equator


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago
When wire is bent in the form of semicircular arc then,l=πr
∴The radius of semicircular arc,r=l/π
Distance between two end points of semicircular wire=2r=π2l​
∴Magnetic moment of semicircular wire
=m×2r=m×π2l​=π2​ml
Butmlis the magnetic moment of straight wire
i.e.,ml=M
∴New magnetic moment=π2​M